Circular No. 6389

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1630–47 F. E. Marshall, Laboratory for High Energy Astrophysics, Goddard

Space Flight Center, reports the detection of the recurrent ultrasoft xray transient 4U 1630–47 in outburst: “4U 1630–47 is thought to be a blackhole x-ray transient system because of its spectrum and the lack of detected pulsations or bursts. A flux in the band 2–10 keV of 0.26 times that of the Crab was observed on Apr. 23.794 UT during a brief scan with the Proportional Counter Array (PCA) onboard the Rossi X-ray Timing Explorer. The flux has increased from 0.004 Crab on Feb. 16.428 and 0.08 Crab on Mar. 12.885. The latest observation occurs at phase 0.29 of the 602-day period for outbursts from this source reported by Parmar et al. (1995, Ap.J. 452, L129). The current outburst is thus unusually delayed from the time expected from previous outbursts. No optical counterpart has been identified for this source, whose position was measured during an earlier observation with the HRI on the Einstein Observatory. A finding chart is given in Parmar et al. (1986, Ap.J. 304, 664). Other blackhole transients also flare at radio wavelengths during x-ray outbursts (e.g., GRS 1915+105, IAUC 6269).”

COMET 29P/SCHWASSMANN-WACHMANN 1 H. Rauer, Observatoire de Paris-Meudon; and H. Böhnhardt, Univer-

sity Observatory, Munich, report observations made at the European Southern Observatory with the Danish 1.54-m telescope (+ DFOSC): “On Apr. 26, comet 29P exhibited an extended coma of elliptical shape (north-south diameter ∼ 3′.5, east-west diameter ∼ 2′). The central coma condensation was starlike (mag R = 17.6). On Apr. 27, this central condensation had R = 14.4, ∼ 3 mag brighter than on the previous day. Despite the suggestion of an outburst, changes in the inner and outer coma are not yet obvious.”

Total visual magnitude estimates (cf. IAUC 6370): Apr. 3.98 UT, 12.1 (J. D. Shanklin, Cambridge, England, 0.30-m refractor); 6.82, 11.2 (A. Baransky, Kiev, Ukraine, 0.11-m reflector); 8.93, 10.7 (J. Carvajal, Madrid, Spain, 0.21-m reflector); 15.24, 11.5 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 16.01, 11.8 (S. Garro, Merlette, France, 0.28-m reflector + CCD); 20.86, 12.2 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector); 22.84, 11.8: (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

1996 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green